AnimeGhoul8863:

The table below shows the distance y, in miles per minute, of a toy car at different times x, in minutes: Time (x) (minutes) 10 20 30 40 Distance (y) (miles) 0.2 0.8 0.1 0.4 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance of the toy car. (4 points) [Choose the value of correlation coefficient from −1, 1, −0.04, 0.99] Part B: What is the value of the slope of the graph of distance versus time between 10 minutes and 20 minutes, and what does the slope represent? (3 points) Part C: Does the data in the table represent correlation or causation? Explain your answer. (3 points)

AnimeGhoul8863:

HERO!! YEW ES HERE PLEASE HELP MEEEEEE T~T

Hero:

Can you screenshot the question please?

Hero:

Okay, so, 1st question is, have you attempted to graph these data points? The best online graph to use for this is http://www.desmos.com

AnimeGhoul8863:

ok what do i do once i have the site i dont know what im suppose to graph

Hero:

You graph the given table of (x,y) points.

AnimeGhoul8863:

ok ill try

AnimeGhoul8863:

AnimeGhoul8863:

^^so it looks like this

Hero:

Okay, so based on that, what is the value of the correlation coefficient? Is the value of the correlation coefficient close to zero or one?

AnimeGhoul8863:

Closer to zero because that makes a straight line

AnimeGhoul8863:

?

Hero:

Correct. So which of the given values for the correlation coefficient is close to zero?

AnimeGhoul8863:

0.1

AnimeGhoul8863:

now what

Hero:

I should ask you, what are the given values of the correlation coefficient they are asking you to choose from?

AnimeGhoul8863:

−1, 1, −0.04, 0.99

AnimeGhoul8863:

its not 1

AnimeGhoul8863:

its not -1

Hero:

Which of those values are closest to zero?

AnimeGhoul8863:

and its not 0.99 so it has to be -0.04 right

Hero:

Correct.

AnimeGhoul8863:

ok so what would be next

Hero:

Next you have to calculate the slope using values for the first two columns of the table.

AnimeGhoul8863:

so 10 and 20 and 0.2 and 0.8?

AnimeGhoul8863:

how would i calculate that? 10 - 20 over 0.2 - 0.8

AnimeGhoul8863:

@Hero

Hero:

You should actually make it so that when you subtract you end up with positive numbers. In other words subtract the smaller numbers from the bigger numbers.

AnimeGhoul8863:

im confused i thought that what i did

AnimeGhoul8863:

so it would be 10-0.2 and 20-0.8?

Hero:

What you've done now is mix up the x's and y's. You subtract y's from y's and x's from x's.

AnimeGhoul8863:

im confused

AnimeGhoul8863:

before i put y-y and x-x but u said u have to do smaller - bigger so i gave you smaller - bigger

AnimeGhoul8863:

can you show me an example please cause i really am confused rn

AnimeGhoul8863:

@Hero

Hero:

All you had to do was swap the order of the y's and x's, not mix them up.

AnimeGhoul8863:

ok

AnimeGhoul8863:

so 20-10 and 0.8-0.2?

Hero:

There you go

AnimeGhoul8863:

ok so its 10 and 0.4

Hero:

Yeah, but you have to simplify that because what you actually have is \(\dfrac{10}{\dfrac{4}{10}}\)

Hero:

\(0.4\) and \(\dfrac{4}{10}\) both mean four-tenths. One is in decimal form the other is in fraction form.

AnimeGhoul8863:

10 over 4 over ten = 25 when simplified

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

Just to let you know there are 3 more questions i need help on after dis if thats okie e.e

Hero:

It's \(10 \div \dfrac{4}{10}\) or \(10\) divided by \(\dfrac{4}{10}\) but yeah 25 is correct.

Hero:

So what does that mean in terms of the slope?

AnimeGhoul8863:

urm it means that the slope is 25?

Hero:

Correct.

AnimeGhoul8863:

so after we find the slope now what

AnimeGhoul8863:

what i have so far Part A: The correlation coefficient to the data is -0.04 Because its the closest to zero.

Hero:

... because the graph of the given data suggests that the correlation coefficient is close to zero.

AnimeGhoul8863:

Ohhh ok that makes more since

Hero:

Whoops, I deleted the wrong thing. Can you re-post that?

AnimeGhoul8863:

This? Part A: The correlation coefficient to the data is -0.04 Because its the closest to zero.

AnimeGhoul8863:

@Hero This is what u wanted?

Hero:

No you posted something else after that. I deleted it accidentally.

AnimeGhoul8863:

Based on the correlation coefficient, describe the relationship between time and distance of the toy car.

AnimeGhoul8863:

Dis?

AnimeGhoul8863:

@Hero x'D yew still there

Hero:

So basically, the correlation coefficient is a way of measuring the strength of relationship between two variables. Values close to one indicate a strong positive relationship. Values close to negative one indicate a strong negative relationship. Values close to zero indicate no relationship. That's your hint.

AnimeGhoul8863:

The relationship between these variables is a No relationship because its closer to zero

AnimeGhoul8863:

^^^If this is correct den were on to Part B:

AnimeGhoul8863:

The slope equals 25. This represents the time it takes for the toy car to hit its maximum?

Hero:

One moment.

Hero:

Do you know the slope formula?

Hero:

If so, then what is it?

AnimeGhoul8863:

y=mx+b?

AnimeGhoul8863:

i thought we already did the slope tho

Hero:

That's the equation of a line, not the slope formula.

AnimeGhoul8863:

den i gots no clue

Hero:

I went with the information you provided because I assumed you knew the slope formula

AnimeGhoul8863:

its been a while

Hero:

\(m = \dfrac{y_2 - y_1}{x_2 - x_1}\). If you don't know it, you should have looked it up.

Hero:

The y values go in the numerator, x values in the denominator.

AnimeGhoul8863:

i did e.e

Hero:

No, you did the reverse of what I just said above.

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AnimeGhoul8863 how would i calculate that? 10 - 20 over 0.2 - 0.8 \(\color{#0cbb34}{\text{End of Quote}}\)

Hero:

That's the proof you put x values over y values

AnimeGhoul8863:

so i use that for the slope?

AnimeGhoul8863:

im so confused

AnimeGhoul8863:

why are we doing the slope all over agian we already figured it out

AnimeGhoul8863:

it equals 25

Hero:

Are you not understanding that the slope is wrong because you put x values over y values when the slope formula suggests to put the y values over the x values?

AnimeGhoul8863:

What are you talking about

AnimeGhoul8863:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AnimeGhoul8863 urm it means that the slope is 25? \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hero Correct. \(\color{#0cbb34}{\text{End of Quote}}\)

Hero:

Yep, and that is wrong for reasons I already mentioned. You'll have to re-calculate it.

AnimeGhoul8863:

can you give me an example of what you mean cause your really not making since rn

Hero:

Example: x = {2, 4} y = {3, 6} Calculate slope Slope formula is \(m = \dfrac{y_2 - y_1}{x_2 - x_1}\) You did the equivalent of this when calculating the slope: \(m = \dfrac{4 - 2}{6 - 3} = \dfrac{2}{3}\) You put the x values in the numerator and the y values in the denominator when in fact the correct slope is this: \(m = \dfrac{6 - 3}{4 - 2} = \dfrac{3}{2}\)

AnimeGhoul8863:

\[\frac{ 20-10 }{ 0.8-0.2 }\]?????

Hero:

That's not correct.

Hero:

Put the y values in the numerator and the x values in the denominator.

AnimeGhoul8863:

\[\frac{ 0.8-0.2 }{ 20-10 }\]????????

Hero:

Finally...

AnimeGhoul8863:

im sorry but this confuses me extremely easily

Hero:

Sorry to hear that. \(\dfrac{\dfrac{6}{10}}{10}\) is what we have now or \(\dfrac{6}{10} \div 10\)

AnimeGhoul8863:

are we looking for exact form or decimal form?

Hero:

I have to go after this so basically \(\dfrac{6}{100}\) is close to a zero slope. In order to make any sense out of this we need to force the denominator to equal one. The only way to do that is to multiply top and bottom by \(\dfrac{1}{100}\)

Hero:

When we do that we get \(\dfrac{\dfrac{6}{100}}{1}\) or \(\dfrac{0.06}{1}\) So the slope means that for every minute the toy car travels, it only travels a distance of six hundredths of a mile.

Hero:

Which makes sense since it is a toy car. We wouldn't expect it to travel far in a minute.

AnimeGhoul8863:

so the slope Equals 0.06 ok that makes since